# Entropy Functions on Two-Dimensional Faces of Polymatroidal Region of Degree Four: Part II: Information Theoretic Constraints Breed New Combinatorial Structures

### v1

@@ II-D Entropy functions on the extreme rays of $\Gamma_{4}$
As we discussed in Subsection II-B, for the V-representation of a face, it can be written as the convex hull of the extreme rays it contains. Therefore, to characterize entropy functions on $2$-dimensional faces of $\Gamma_{4}$, we need first to characterize the entropy functions on the extreme rays of $\Gamma_{4}$. For the eleven types of extreme rays, entropy functions on seven of them are matroidal and have been characterized in [6], while those containing $\hat{U}^{i}_{2,5}$ or $\hat{U}^{i}_{3,5}$ are characterized in Part I of this series [22], and those containing $V^{\alpha}_{8}$ are non-entropic. Theorems characterize these entropy functions are listed below without proofs.
@@
([6, Proposition 2]) For $E=U_{2,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log k$ for positive integer $k\neq 2,6$.
@@
([22, Theorem II-C.4]) For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{2,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log k$ for integer $k>0$.
@@
([22, Theorem II-C.5]) For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{3,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log k$ for integer $k>0$.
@@
Remark In Part I [22], Theorem II-D.4(resp. II-D.5) on extreme ray containing $\hat{U}^{i}_{2,5}$(resp. $\hat{U}^{i}_{3,5}$) was proved by the construction of a specific $\mathrm{MVOA}(\hat{U}^{i}_{2,5},v)$(resp. $\mathrm{MVOA}(\hat{U}^{i}_{3,5},v)$) for all $v\geq 1$. However, the correspondence between a pair of Latin squares of the first kind and the second kind and an $\mathrm{MVOA}(\hat{U}^{i}_{2,5},v)$(resp. a Latin cube of the second kind and an $\mathrm{MVOA}(\hat{U}^{i}_{3,5},v)$) setting up in Subsection II-C provides a general construction of the problem. Such construction also sheds some lights on the open question [8, Question 1].
@@
III Characterization of entropy functions on all Two-dimensional faces of $\Gamma_{4}$
In this section, we characterize the entropy functions on $2$-dimensional faces of $\Gamma_{4}$. We embed each face $F=(E_{1},E_{2})$ in the first octant of a $2$-dimensional cartesian coordinate system whose axis are labeled by $a$ and $b$. Thus, for each $(a,b)$, $a,b\geq 0$, it represents the polymatroid $a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{i},i=1,2$ is the rank function of the minimal integer polymatroid in $E_{i}$, respectively. Throughtout this paper, for a random vector $(X_{i},i\in N_{4})$ or its subvectors, we assume each $X_{i}$ is distributed on a finite set $\mathcal{X}_{i}$, and for each $x_{i}\in X_{i}$, $p(x_{i})>0$.
@@ III Characterization of entropy functions on all Two-dimensional faces of $\Gamma_{4}$
([22, Lemma 1]) If $X_{1}$ and $X_{2}$ are independent and for any $p(x_{1},x_{2},x_{3})>0$, $p(x_{1})=p(x_{2})$, then $X_{1}$ and $X_{2}$ are uniformly distributed on $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, respectively, $|\mathcal{X}_{1}|=|\mathcal{X}_{2}|$ and $H(X_{1})=H(X_{2})$.
@@ III-A Entropy functions on faces with extreme rays containing $U_{2,4}$ and one rank-1 matroid
by [9, Theorem 1], which is an outer bound on the entropy region of $F$.
@@
Figure 2: The entropy region on $F$
@@
$H(X_{1})$ $=-\sum_{j=1}^{t_{1}}n^{(j)^{\prime}}\dfrac{p^{\prime}_{j}}{n^{(j)^{\prime}}}\log\dfrac{p^{\prime}_{j}}{n^{(j)^{\prime}}}$ (III.1.22)
$=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}},$ (III.1.23)
@@
$H(X_{1})$ $=H(X_{2})=H(X_{3})$ (III.1.24)
$=a+b,$ (III.1.25)
$H(X_{4})$ $=a.$ (III.1.26)
By (III.1.23) and $n^{(j)}\geq v$,
$a+b$ $=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}}$ (III.1.27)
$\geq H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log v$ (III.1.28)
$\geq\log v.$ (III.1.29)
@@
$n_{3}^{(j)}=\dfrac{p^{\prime}_{j}}{p(x_{3})}=\dfrac{p_{j}}{p(x_{1})}=n^{(j)},$ (III.1.31)
@@
$n^{(j)^{\prime}}=\sum_{i\in A_{j}}n^{(i)}\geq v+1.$ (III.1.32)
By (III.1.23) and (III.1.25),
$a+b=H(X_{1})$ $=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}}$ (III.1.33)
$\geq H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{(v+1)}$ (III.1.34)
$\geq\log{(v+1)},$ (III.1.35)
@@
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (III.1.36)
$=p(x_{1},x_{2},x_{4})$ (III.1.37)
$=p(x_{1},x_{3},x_{4})$ (III.1.38)
$=p(x_{2},x_{3},x_{4}),$ (III.1.39)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (III.1.40)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (III.1.41)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.1.42)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.1.43)
$p(x_{3},x_{4})$ $=p(x_{4})p(x_{4}),$ (III.1.44)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1},x_{2},x_{3},x_{4}),$ (III.1.45)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1},x_{2},x_{3},x_{4}),$ (III.1.46)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1},x_{2},x_{3},x_{4}).$ (III.1.47)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1},x_{2},x_{3},x_{4}),$ (III.1.48)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1},x_{2},x_{3},x_{4}).$ (III.1.49)
According to (III.1.36), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (III.1.45), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{2}).$ (III.1.50)
Together with (III.1.37), we obtain
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2}).$ (III.1.51)
@@
$b\geq\log 2.$ (III.1.56)
Consider the bipartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}$ and $(x_{1},x_{2})\in E$ if and only if $p(x_{1},x_{2})>0$. Assume $G$ has $t$ connected components and $|\mathcal{X}_{3}|=|\mathcal{X}_{4}|=v$ for $v=2,6$. By (III.1.36), (III.1.37) and (III.1.50), we have $p(x_{1},x_{2})=p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})$, which implies that $X_{i}$ is a function of $X_{1}$ and $X_{2}$, $i=3,4$. Then each edge $(x_{1},x_{2})$ can be colored by a unique $x_{3}\in\mathcal{X}_{3}$ and a unique $x_{4}\in\mathcal{X}_{4}$. By (III.1.36) and (III.1.52), we have $p(x_{1},x_{3})=p(x_{1},x_{2},x_{3})$. Thus any two edges incident to $x_{1}$ are colored by different $x_{3}\in\mathcal{X}_{3}$. Since $X_{1}$ and $X_{3}$ are independent, all colors $x_{3}\in\mathcal{X}_{3}$ will occur at least once on the edges adjacent to $x_{1}$. Hence, each $x_{1}$ is incident to $k$ edges. It holds for each $x_{2}$ as well by symmetry. We denote the number of the vertices in $\mathcal{X}_{i}$ by $n_{i}^{(j)},i=1,2,j=1,2,\cdots,t$ in the connected component $C_{j}$ and the probability mass of $C_{j}$ by $p_{j}$. In each connected component $C_{j}$, the number of edges is $n_{1}^{(j)}k=n_{2}^{(j)}k$, which implies $n_{1}^{(j)}=n_{2}^{(j)}$ and so it can be simplified to $n^{(j)}$. As $p(x_{1},x_{2})=\sum_{x_{3},x_{4}}p(x_{1},x_{2},x_{3},x_{4})$, there exist $x_{3},x_{4}$ such that $p(x_{1},x_{2},x_{3},x_{4})>0$. Due to (III.1.52) and (III.1.53), we have $p(x_{1},x_{3})=p(x_{2},x_{3})$. By (III.1.40) and (III.1.42), replacing $p(x_{1},x_{3})$ and $p(x_{2},x_{3})$ by $p(x_{1})p(x_{3})$ and $p(x_{2})p(x_{3})$, we obtain
$p(x_{1})=p(x_{2}),$ (III.1.57)
@@
$p(x_{1})=p(x_{2})=\dfrac{p_{j}}{n^{(j)}}.$ (III.1.58)
By (III.1.40), (III.1.51) and (III.1.52), the probability mass of the edges are equal to
$p(x_{1},x_{2})$ $=p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{3})=p(x_{1})p(x_{3}).$ (III.1.59)
Note that $X_{3}$ is uniformly distributed on $\mathcal{X}_{3}$ and $|\mathcal{X}_{3}|=v$. Replacing $p(x_{3})$ by $\frac{1}{v}$ in (III.1.59), we have
$p(x_{1},x_{2})=p(x_{1})p(x_{3})=\dfrac{p(x_{1})}{v}=\dfrac{p_{j}}{n^{(j)}v}.$ (III.1.60)
@@
$H(X_{1})$ $=-\sum_{j=1}^{t}n^{(j)}\dfrac{p_{j}}{n^{(j)}}\log\dfrac{p_{j}}{n^{(j)}}$ (III.1.61)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}.$ (III.1.62)
$H(X_{1},X_{2})$ $=-\sum_{j=1}^{t}n^{(j)}v\dfrac{p_{j}}{n^{(j)}v}\log\dfrac{p_{j}}{n^{(j)}v}$ (III.1.63)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}v$ (III.1.64)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}+\sum_{j=1}^{t}p_{j}\log v$ (III.1.65)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}+\log v.$ (III.1.66)
@@
$H(X_{1})$ $=a+b,$ (III.1.67)
$H(X_{1},X_{2})$ $=2a+b.$ (III.1.68)
Equating the above equations with (III.1.62) and (III.1.66), we obtain
$a$ $=\log v,$ (III.1.69)
$b$ $=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}-\log v.$ (III.1.70)
Now we show that for each connected component, there exists an $(x_{3},x_{4})$ that occurs at least twice. By (III.1.44), $X_{3}$ and $X_{4}$ are independent. Since $|\mathcal{X}_{3}|=|\mathcal{X}_{4}|=v$, the number of the pairs $(x_{3},x_{4})$ is $v^{2}$. The probability mass of each $(x_{3},x_{4})$ is
$p(x_{3},x_{4})=p(x_{3})p(x_{4})=\dfrac{1}{v^{2}}.$ (III.1.71)
@@
$p(x_{3},x_{4})$ $=\sum_{x_{1},x_{2}}p(x_{1},x_{2},x_{3},x_{4})\geq\dfrac{2p_{j}}{n^{(j)}v}.$ (III.1.74)
Together with (III.1.71), we obtain
$\dfrac{1}{v^{2}}\geq\dfrac{2p_{j}}{n^{(j)}v},$ (III.1.75)
@@
To prove the theorem, it remains to show that all $(a,b)\in F$ are entropic if $a=\log 2$ or $\log 6$, $b\geq 6$, which can be implied by the fact that polymatroids $\mathbf{h}=(\log 2,\log 2)$ and $(\log 6,\log 2)$ are entropic. For $v=2,6$, let $\mathbf{T}_{v}$ be a $2v^{2}\times 4$ array with entries in $\mathbb{I}_{2v}$ such that
@@
$\mathbf{T}_{6}\ =\ \begin{matrix}0&0&0&0\\ 0&1&1&1\\ 0&2&2&2\\ 0&3&3&3\\ 0&4&4&4\\ 0&5&5&5\\ 1&0&5&1\\ 1&1&0&2\\ 1&2&1&3\\ 1&3&2&4\\ 1&4&3&5\\ 1&5&4&0\\ 2&0&4&5\\ 2&1&5&0\\ 2&2&0&1\\ 2&3&1&2\\ 2&4&2&3\\ 2&5&3&4\\ 3&0&3&3\\ 3&1&4&4\\ 3&2&5&5\\ 3&3&0&0\\ 3&4&1&1\\ 3&5&2&2\\ 4&0&2&4\\ 4&1&3&5\\ 4&2&4&0\\ 4&3&5&1\\ 4&4&0&2\\ 4&5&1&3\\ 5&0&1&2\\ 5&1&2&3\\ 5&2&3&4\\ 5&3&4&5\\ 5&4&5&0\\ 5&5&0&1\\ \end{matrix}\qquad\qquad\begin{matrix}6&6&0&3\\ 6&7&1&4\\ 6&8&2&5\\ 6&9&3&0\\ 6&10&4&1\\ 6&11&5&2\\ 7&6&5&4\\ 7&7&0&5\\ 7&8&1&0\\ 7&9&2&1\\ 7&10&3&2\\ 7&11&4&3\\ 8&6&4&2\\ 8&7&5&3\\ 8&8&0&4\\ 8&9&1&5\\ 8&10&2&0\\ 8&11&3&1\\ 9&6&3&0\\ 9&7&4&1\\ 9&8&5&2\\ 9&9&0&3\\ 9&10&1&4\\ 9&11&2&5\\ 10&6&2&1\\ 10&7&3&2\\ 10&8&4&3\\ 10&9&5&4\\ 10&10&0&5\\ 10&11&1&0\\ 11&6&1&5\\ 11&7&2&0\\ 11&8&3&1\\ 11&9&4&2\\ 11&10&5&3\\ 11&11&0&4\end{matrix}$
@@
It can be checked that the entropy function is in $F$. Then by (III.1.67), (III.1.68) and (III.1.82), we have $a=\log v$ and $b=\log 2$. The proof is accomplished. ∎
@@
$p_{X_{1},X_{4}}(1,x_{4})=p_{X_{1}}(1)p_{X_{4}}(x_{4})>0,$ (III.1.102)
which implies that either $p_{X_{1},X_{2},X_{3},X_{4}}(1,0,1,x_{4})>0$ or $p_{X_{1},X_{2},X_{3},X_{4}}(1,1,0,x_{4})>0$. Since $p_{X_{1}X_{2}X_{3}X_{4}}(x_{1},x_{2},x_{3},x_{4})=p_{X_{2}X_{4}}(x_{2},x_{4})$ by (III.1.100),

### v2

@@ II-D Entropy functions on the extreme rays of $\Gamma_{4}$
As we discussed in Subsection II-B, for the V-representation of a face, it can be written as the convex hull of the extreme rays it contains. Therefore, to characterize entropy functions on $2$-dimensional faces of $\Gamma_{4}$, we need first to characterize the entropy functions on the extreme rays of $\Gamma_{4}$. For the eleven types of extreme rays, the entropy functions on seven of them are matroidal and have been characterized in [7], while those containing $\hat{U}^{i}_{2,5}$ or $\hat{U}^{i}_{3,5}$ are characterized in Part I of this series [23], and those containing $V^{\alpha}_{8}$ are non-entropic. Theorems that characterize these entropy functions are listed below without proof.
@@
([7, Proposition 2]) For $E=U_{2,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log k$ for positive integer $k\neq 2,6$.
@@
([23, Theorem II-C.4]) For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{2,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log k$ for integer $k>0$.
@@
([23, Theorem II-C.5]) For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{3,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log k$ for integer $k>0$.
@@
Remark In Part I [23], Theorem II-D.4(resp. II-D.5) on extreme ray containing $\hat{U}^{i}_{2,5}$(resp. $\hat{U}^{i}_{3,5}$) was proved by the construction of a specific $\mathrm{MVOA}(\hat{U}^{i}_{2,5},v)$(resp. $\mathrm{MVOA}(\hat{U}^{i}_{3,5},v)$) for all $v\geq 1$. However, the correspondence between a pair of Latin squares of the first kind and the second kind and an $\mathrm{MVOA}(\hat{U}^{i}_{2,5},v)$(resp. a Latin cube of the second kind and an $\mathrm{MVOA}(\hat{U}^{i}_{3,5},v)$) setting up in Subsection II-C provides a general construction of the problem. Such construction also sheds some light on the open question [9, Question 1].
@@
III Characterization of entropy functions on Two-dimensional faces of $\Gamma_{4}$
In this section, we characterize the entropy functions on $2$-dimensional faces of $\Gamma_{4}$. We embed each face $F=(E_{1},E_{2})$ in the first octant of a $2$-dimensional Cartesian coordinate system whose axes are labeled by $a$ and $b$. Thus, for each $(a,b)$, $a,b\geq 0$, it represents the polymatroid $a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{i},i=1,2$, is the rank function of the minimal integer polymatroid in $E_{i}$, respectively. Throughout this paper, for a random vector $(X_{i},i\in N_{4})$ or its subvectors, we assume each $X_{i}$ is distributed on a finite set $\mathcal{X}_{i}$, and for each $x_{i}\in X_{i}$, $p(x_{i})>0$.
@@ III Characterization of entropy functions on Two-dimensional faces of $\Gamma_{4}$
([23, Lemma 1]) If $X_{1}$ and $X_{2}$ are independent and for any $p(x_{1},x_{2},x_{3})>0$, $p(x_{1})=p(x_{2})$, then $X_{1}$ and $X_{2}$ are uniformly distributed on $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, respectively, $|\mathcal{X}_{1}|=|\mathcal{X}_{2}|$ and $H(X_{1})=H(X_{2})$.
@@ III-A Entropy functions on faces with extreme rays containing $U_{2,4}$ and one rank-1 matroid
by [10, Theorem 1], which is an outer bound on the entropy region of $F$.
@@
Figure 2: Inner and outer bounds on $F^{*}$
@@
$H(X_{1})$ $=-\sum_{j=1}^{t_{1}}n^{(j)^{\prime}}\dfrac{p^{\prime}_{j}}{n^{(j)^{\prime}}}\log\dfrac{p^{\prime}_{j}}{n^{(j)^{\prime}}}$ (III.1.23)
$=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}},$ (III.1.24)
@@
$H(X_{1})$ $=H(X_{2})=H(X_{3})$ (III.1.25)
$=a+b,$ (III.1.26)
$H(X_{4})$ $=a.$ (III.1.27)
By (III.1.24) and $n^{(j)}\geq v$,
$a+b$ $=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}}$ (III.1.28)
$\geq H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log v$ (III.1.29)
$\geq\log v.$ (III.1.30)
@@
$n_{3}^{(j)}=\dfrac{p^{\prime}_{j}}{p(x_{3})}=\dfrac{p_{j}}{p(x_{1})}=n^{(j)},$ (III.1.32)
@@
$n^{(j)^{\prime}}=\sum_{i\in A_{j}}n^{(i)}\geq v+1.$ (III.1.33)
By (III.1.24) and (III.1.26),
$a+b$ $=H(X_{1})=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}}$ (III.1.34)
$\geq H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{(v+1)}$ (III.1.35)
$\geq\log{(v+1)},$ (III.1.36)
@@
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.1.37)
$=p(x_{1},x_{2},x_{4})$ (III.1.38)
$=p(x_{1},x_{3},x_{4})$ (III.1.39)
$=p(x_{2},x_{3},x_{4}),$ (III.1.40)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (III.1.41)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (III.1.42)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.1.43)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.1.44)
$p(x_{3},x_{4})$ $=p(x_{4})p(x_{4}),$ (III.1.45)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (III.1.46)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.1.47)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (III.1.48)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.1.49)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}).$ (III.1.50)
According to (III.1.37), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (III.1.46), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{2}).$ (III.1.51)
Together with (III.1.38), we obtain
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2}).$ (III.1.52)
@@
$b\geq\log 2.$ (III.1.57)
Consider the bipartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}$ and $(x_{1},x_{2})\in E$ if and only if $p(x_{1},x_{2})>0$. Assume $G$ has $t$ connected components and $|\mathcal{X}_{3}|=|\mathcal{X}_{4}|=v$ for $v=2,6$. By (III.1.37), (III.1.38) and (III.1.51), we have $p(x_{1},x_{2})=p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})$, which implies that $X_{i}$ is a function of $X_{1}$ and $X_{2}$, $i=3,4$. Then each edge $(x_{1},x_{2})$ can be colored by a unique $x_{3}\in\mathcal{X}_{3}$ and a unique $x_{4}\in\mathcal{X}_{4}$. By (III.1.37) and (III.1.53), we have $p(x_{1},x_{3})=p(x_{1},x_{2},x_{3})$. Thus, any two edges incident to $x_{1}$ are colored by different $x_{3}\in\mathcal{X}_{3}$. Since $X_{1}$ and $X_{3}$ are independent, all colors $x_{3}\in\mathcal{X}_{3}$ will occur at least once on the edges adjacent to $x_{1}$. Hence, each $x_{1}$ is incident to $k$ edges. It holds for each $x_{2}$ as well by symmetry. We denote the number of the vertices in $\mathcal{X}_{i}$ by $n_{i}^{(j)},i=1,2,j=1,2,\cdots,t$ in the connected component $C_{j}$ and the probability mass of $C_{j}$ by $p_{j}$. In each connected component $C_{j}$, the number of edges is $n_{1}^{(j)}k=n_{2}^{(j)}k$, which implies $n_{1}^{(j)}=n_{2}^{(j)}$ and so it can be simplified to $n^{(j)}$. As $p(x_{1},x_{2})=\sum_{x_{3},x_{4}}p(x_{1},x_{2},x_{3},x_{4})$, there exist $x_{3},x_{4}$ such that $p(x_{1},x_{2},x_{3},x_{4})>0$. Due to (III.1.53) and (III.1.54), we have $p(x_{1},x_{3})=p(x_{2},x_{3})$. By (III.1.41) and (III.1.43), replacing $p(x_{1},x_{3})$ and $p(x_{2},x_{3})$ by $p(x_{1})p(x_{3})$ and $p(x_{2})p(x_{3})$, we obtain
$p(x_{1})=p(x_{2}),$ (III.1.58)
@@
$p(x_{1})=p(x_{2})=\dfrac{p_{j}}{n^{(j)}}.$ (III.1.59)
By (III.1.41), (III.1.52) and (III.1.53), the probability mass of the edges are equal to
$p(x_{12})$ $=p(x_{1234})=p(x_{1},x_{3})=p(x_{1})p(x_{3}).$ (III.1.60)
Note that $X_{3}$ is uniformly distributed on $\mathcal{X}_{3}$ and $|\mathcal{X}_{3}|=v$. Replacing $p(x_{3})$ by $\frac{1}{v}$ in (III.1.60), we have
$p(x_{1},x_{2})=p(x_{1})p(x_{3})=\dfrac{p(x_{1})}{v}=\dfrac{p_{j}}{n^{(j)}v}.$ (III.1.61)
@@
$H(X_{1})$ $=-\sum_{j=1}^{t}n^{(j)}\dfrac{p_{j}}{n^{(j)}}\log\dfrac{p_{j}}{n^{(j)}}$ (III.1.62)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}.$ (III.1.63)
$H(X_{1},X_{2})=-\sum_{j=1}^{t}n^{(j)}v\dfrac{p_{j}}{n^{(j)}v}\log\dfrac{p_{j}}{n^{(j)}v}$ (III.1.64)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}v$ (III.1.65)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}+\sum_{j=1}^{t}p_{j}\log v$ (III.1.66)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}+\log v.$ (III.1.67)
@@
$H(X_{1})$ $=a+b,$ (III.1.68)
$H(X_{1},X_{2})$ $=2a+b.$ (III.1.69)
Equating the above equations with (III.1.63) and (III.1.67), we obtain
$a$ $=\log v,$ (III.1.70)
$b$ $=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}-\log v.$ (III.1.71)
Now we show that for each connected component, there exists an $(x_{3},x_{4})$ that occurs at least twice. By (III.1.45), $X_{3}$ and $X_{4}$ are independent. Since $|\mathcal{X}_{3}|=|\mathcal{X}_{4}|=v$, the number of the pairs $(x_{3},x_{4})$ is $v^{2}$. The probability mass of each $(x_{3},x_{4})$ is
$p(x_{3},x_{4})=p(x_{3})p(x_{4})=\dfrac{1}{v^{2}}.$ (III.1.72)
@@
$p(x_{3},x_{4})$ $=\sum_{x_{1},x_{2}}p(x_{1},x_{2},x_{3},x_{4})\geq\dfrac{2p_{j}}{n^{(j)}v}.$ (III.1.75)
Together with (III.1.72), we obtain
$\dfrac{1}{v^{2}}\geq\dfrac{2p_{j}}{n^{(j)}v},$ (III.1.76)
@@
To prove the theorem, it remains to show that all $(a,b)\in F$ are entropic if $a=\log 2$ or $\log 6$, $b\geq\log 2$, which can be implied by the fact that polymatroids $\mathbf{h}=(\log 2,\log 2)$ and $(\log 6,\log 2)$ are entropic. For $v=2,6$, let $\mathbf{T}_{v}$ be a $2v^{2}\times 4$ array with entries in $\mathbb{I}_{2v}$ such that
@@
[…]
@@
It can be checked that the entropy function is in $F$. Then by (III.1.68), (III.1.69) and (III.1.82), we have $a=\log v$ and $b=\log 2$. The proof is accomplished. ∎
@@
$p_{X_{1}X_{4}}(1,x_{4})=p_{X_{1}}(1)p_{X_{4}}(x_{4})>0,$ (III.1.102)
which implies that either $p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4})>0$ or $p_{X_{1}X_{2}X_{3}X_{4}}(1,1,0,x_{4})>0$. Since $p_{X_{1}X_{2}X_{3}X_{4}}(x_{1},x_{2},x_{3},x_{4})=p_{X_{2}X_{4}}(x_{2},x_{4})$ by (III.1.100),